# A Review of Large Language Models and Autonomous Agents in Chemistry

### v1

@@ 4 LLM-based autonomous agents
The term “agent” originates in philosophy, referring to entities capable of making decisions347. Hence, in artificial intelligence, an “agent” is a system that can perceive its environment, make decisions, and act upon them in response to external stimuli348. Language has enabled humans to decide and act to make progress in response to the environment and its stimuli, and so LLMs are naturally ideal for serving as the core of autonomous agents. Thus, in agreement with Gao et al. 349, we define a “language agent” as a model or program (typically based on LLMs) that receives an observation from its environment and executes an action in this environment. Here, environment means a set of tools and a task. Hence, “LLM-based autonomous agents” refer to language agents whose core is based on an LLM model. Comprehensive analyses of these agents are available in the literature350, 348, 349, but this section highlights key aspects to prepare the reader for future discussions.
There is no agreed definition of the nomenclature to be used to discuss agents. For instance, Gao et al. 349 created a classification scheme that aims to group agents by their autonomy in biological research. This means a level 0 agent has no autonomy and can only be used as a tool, while a level 3 agent can independently create hypotheses, design experiments, and reason. Another previously suggested framework highlights agents’ construction, application, and evaluation as a parallel between agents and neural networks. Following this perspective, Wang et al. 350 categorizes agent components into four modules: profiling, memory, planning, and action. In contrast, Weng 351 also identifies four elements — memory, planning, action, and tools — but with a different emphasis. Meanwhile, Xi et al. 348 proposes a division into three components: brain, perception, and action, integrating profiling, memory, and planning within the brain component, where the brain is typically an LLM. Recently, Sumers et al. 352 proposed Cognitive Architectures for Language Agents (CoALA), a conceptual framework to generalize and ease the design of general-purpose cognitive language agents. In their framework, a larger cognitive architecture composed of modules and processes is defined. CoALA defines a memory, decision-making, and core processing module, in addition to an action space composed of both internal and external tools. While internal tools mainly interact with the memory to support decision-making, external tools compose the environment, as illustrated in Figure 5. Given a task that initiates the environment, the “decision process” runs continuously in a loop, receiving observations and executing actions until the task is completed. For more details, read Reference [352].
@@
In this review, we define autonomous agent systems as models (typically LLMs) that continuously receive observations and execute actions within an environment, as described by Gao et al. 349. Nevertheless, in contrast to CoALA349, we will rename “internal tools” as “agent modules” and “external tools” simply as “tools”, for clarity. The agent consists of trainable decision-making components such as the LLM itself, policy, memory, and reasoning scheme. In contrast, the environment comprises non-trainable elements like the task to be completed, Application Programming Interface (API) access, interfaces with self-driving labs, dataset access, and execution of external code. By referring to decision-making components as agent modules, we emphasize their inclusion as parts of the agent. By referring to non-trainable elements as tools, we highlight their role as part of the environment. We discuss five main types of actions. As shown in Figure 5, three of the five, memory, planning and reasoning, and profiling are agent modules. The remaining two actions, tools and perception, are part of the environment. Since the perception is how the agent interacts with the environment and is not a trainable decision, we therefore included it as part of the environment.
@@ 4.1 Memory module
The role of the memory module is to store and recall information from past interactions and experiences to inform future decisions and actions. There are multiple types of memory in agents, namely sensory memory, short-term memory, and long-term memory. A major challenge in using agents is the limited context window, which restricts the amount of in-context information and can lead to information loss, thereby impacting the effectiveness of short-term and long-term memory. Solutions involve summarizing memory content353, compressing memories into vectors354, 355, 356, and utilizing vector databases357 or combinations thereof358, with various databases available such as ChromaDB, FAISS, Pinecone, Weaviate, Annoy, and ScaNN.359 Addressing these challenges to enhance agent memory continues to be a significant area of research360. Sensory, or procedural, memory is knowledge embedded into the model’s parameters during pretraining and/or in heuristics implemented into the agent’s code. Short-term, or working, memory includes the agent’s finite knowledge during a task, incorporating interaction history and techniques like in-context learning361 (ICL), which leverages the limited input’s context length for information retention. Long-term memory involves storing information externally, typically through an embedded vector representation in an external database. In the original CoALA349 paper, long-term memory is further divided into two different types of long-term memory: episodic, which registers previous experiences, and semantic, which stores general information about the world.
@@ 4.2 Planning and reasoning modules
The planning and reasoning module is made of two components. Planning involves identifying a sequence of actions required to achieve a specified goal. In the context of language agents, this means generating steps or strategies that the model can follow to solve a problem or answer a question, which can be enhanced with retrieval from previous experiences,362 and from feedback from post-execution reasoning363, 364. Reasoning refers to the process of drawing conclusions or making decisions based on available information and logical steps. For example, there are studies that demonstrate the benefits of LLM reasoning for question answering, where new context tokens can be integrated in a step-by-step way to guide the model towards more accurate answers365, 366, 367, 368, 369, 370. One popular reasoning strategy is Chain-of-Thought (CoT),367, 371, 372, 91, 373, 374 a reasoning strategy which substantially boosts QA performance by generating intermediate reasoning steps. CoT thereby reduces hallucinations and enhances interpretability, as demonstrated by improved results in models like PaLM375 and GPT-3 with benchmarks like GSM8K376, SVAMPs377, and MAWPS378.
@@
In advanced reasoning, final tasks are often decomposed into intermediary ones using a cascading approach, similar to Zero-shot-CoT366 and RePrompt363. However, while CoT is considered as single-path reasoning, CoT extensions like Tree-of-Thoughts369, Graph-of-Thoughts379, Self-consistent CoT368, and Algorithm-of-Thoughts380 offer multi-path reasoning. Furthermore, other models have pitted multiple agents against each other to debate or discuss various reasoning paths,381, 382, 383 while others use external planners to create plans384, 385. A feedback step during the execution of the plan was a further extension of the CoT ideas; this enables agents to refine their actions based on environmental responses adaptively, which is crucial for complex tasks.386, 387
@@
Another interesting reasoning scheme is the Chain-of-Verification(CoVe), 364 where once an answer is generated, another LLM is prompted to generate a set of verification questions to check for agreement between the original answer and the answers to the verification questions such that the final answer can be refined. The ReAct369 – Reason+Act – model proposes adding an observation step after acting. This means the LLM first reasons about the task and determines the necessary step for its execution, it performs the action and then observes the action’s result. Reasoning on that result, it can subsequently perform the following step. Similarly, Reflexion91 also implements a reasoning step after executing an action. However, Reflexion implements an evaluator and self-reflection LLMs to not only reason about each step but also to evaluate the current trajectory the agent is following using a long-term memory module. As the context increases, it may become challenging for agents to deal with the long prompt. Aiming to solve this issue, the Chain-of-Agents (CoA)388 extends reasoning schemes that leverage multi-agent collaboration to reason over long contexts. This framework employs workers and manager agents to process and synthesize information to generate the final response. CoA demonstrated improvements of up to 10% when compared against an RAG baseline.
@@
ReAct and Reflexion are closed-ended approaches where the agent starts with all the tools and must determine which to use. To address more open-world challenges, Wang et al. 389 introduced the Describe, Explain, Plan, and Select (DEPS) method, which extends this approach. Lastly, human inputs can also be used to provide feedback to the agent. Providing feedback using a human-in-the-loop approach is particularly interesting in fields where safety is a main concern.
@@ 4.3 Profiling module
LLMs can be configured to perform in specific roles, such as coders, professors, students, and domain experts, through a process known as profiling. Language agents can thus incorporate the profile through the LLM or through the agent code. The profiling approach involves inputting psychological characteristics to the agent, significantly impacting its decision-making process 390, 391, 392, 393. Profiling enables the creation of multi-agent systems that simulate societal interactions, with each agent embodying a unique persona within the group 362, 394. The most prevalent technique for profiling, called “handcrafting”, requires manually defining the agent’s profile, often through prompts or system messages 395, 396. While profiling can also be automated with LLMs 397, that automation method may only be suited for generating large numbers of agents since it offers less control over their overall behavior. An interesting application of profiling is the development of agent sets that reflect demographic distributions 398.
@@ 4.4 Perception
Perception is an analog to the human sensory system, which interprets multimodal information such as text, images, or auditory data, transforming it into a format comprehensible by LLMs, as demonstrated by SAM399, GPT4-V400, LLaVa401, Fuyu8B109, and BuboGPT402. In our proposed architecture, the perception is responsible for converting the task and the observations to a data representation that can be understood by the agent. Moreover, advancements in LLMs have led to the development of even more versatile models, such as the any-to-any Next-GPT110 and the any-to-text Macaw-LLM403. Employing multimodal LLMs, which are introduced above, in decision-making processes can simplify perception tasks for agents, with several studies exploring their use in autonomous systems404, 405.
@@ 4.5 Tools
The decision process is composed of three main steps: proposal, evaluation, and selection. During the proposal, one or more action candidates are selected using reasoning369, code structures406, 362, or simply by selecting every tool available407.365, 408, 368 The evaluation process consists of evaluating each selected action according to some metric to predict which action would bring more value to the agent. Lastly, the action is selected and executed.
Given that pretrained parameters (sensory memory) are limited, the model must use tools for complex tasks in order to provide reliable answers. However, LLMs need to learn how to interact with the action space and how and when to use those tools most accurately.409 LLMs can be pretrained or fine-tuned with examples of tool use, enabling them to operate tools and directly retrieve tool calls from sensory memory during a zero-shot generation.410 Recent studies investigate this approach, particularly focusing on open-source LLMs.411, 412, 413
As foundational AI models become more advanced, their abilities can be expanded. It was shown that general-purpose foundation models can reason and select tools even with no fine-tuning. For example, MRKL414 (pronounced “miracle”) implements an extendable set of specialized tools known as neuro-symbolic modules and a smart “router” system to retrieve the best module based on the textual input. These neuro-symbolic modules are designed to handle specific tasks or types of information, equipped with built-in capabilities and task-relevant knowledge. This pre-specialization allows the model to perform domain-specific tasks without needing a separate, domain-specific dataset. This design addresses the problem of LLMs lacking domain-specific knowledge and eliminates the need for the costly and time-consuming LLM fine-tuning step, using specialized data annotation.415 The router can receive support from a reasoning strategy to help select the tools415 or follow a previously created plan389. Recent advances have shown that LLMs can develop new tools of their own416, 417, 418, enabling agents to operate, as needed, in dynamic and unpredictable “open-worlds”, on unseen problems as illustrated by Voyager406. This capability allows agents to evolve and improve continually.
@@ 5 LLM-Based autonomous agents in scientific research
Table 4: Scientific LLM systems and agents. We identify the studies we classified as an agent with the icon [Uncaptioned image]419 and multi-agent systems with the icon [Uncaptioned image][Uncaptioned image]. [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]420 mean the agent bases his behavior on sensory, short, and long memory components, respectively. Besides the textual capabilities of LLM-based agents, [Uncaptioned image]421 and [Uncaptioned image]422 mean the agent has additional audio and visual perception, respectively.
@@
LLaMP423[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools for database access, literature search, and atomistic simulations 2024.06
@@
CRISPR-GPT424[Uncaptioned image][Uncaptioned image] [Uncaptioned image] Tool for gene editing experiments design 2024.04
@@
TAIS425[Uncaptioned image][Uncaptioned image] [Uncaptioned image] Tools for gene expression data analysis 2024.02
SciAgent426[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Trained Mistral for tool usage. Evaluated it using MathToolBench’s tools 2024.02
STORM427 [Uncaptioned image][Uncaptioned image] Article writing using retrieval from multi-LLM conversations and pre-generated outline 2024.02
Völker et al. 428 [Uncaptioned image] Regression with ICL and text retrieval 2024.02
ProtAgent429 [Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools for proteins information retrieval, analyzing, de novo design, and 3D folded structure generation 2024.01
PaperQA430[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools to search the scientific literature, gather evidence, and answer questions 2023.12
WikiCrow431[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Uses PaperQA as a tool 2023.12
Coscientist46[Uncaptioned image] [Uncaptioned image] Tools for running Python code, web-searching, and interacting with lab equipment 2023.12
Eunomia432[Uncaptioned image] [Uncaptioned image] Tools for literature and dataset searching and a chain-of–verification loop 2023.12
CALMS433[Uncaptioned image] [Uncaptioned image] Tools for using the Materials Project API, designing experiments, and using a hardware API to perform the experiment 2023.12
CoQuest434[Uncaptioned image] [Uncaptioned image] Research question generations and tools for literature visualization using a graph organization 2023.10
eXpertAI286[Uncaptioned image] [Uncaptioned image] Tools for applying XAI methods 2023.11
BioPlanner435[Uncaptioned image] [Uncaptioned image] Tool for protocol searching in the BioProt dataset 2023.10
IBM ChatChem436[Uncaptioned image] [Uncaptioned image] Tools for cheminformatics and accessing GT4SD and HuggingFace models 2023.09
@@
ChatMOF437[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools for database search, property prediction, and MOF’s structure generation 2023.08
@@
AmadeusGPT45[Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools for writing and executing code for computer vision, machine learning, and spatial-temporal reasoning 2023.07
@@
i-Digest438 [Uncaptioned image] [Uncaptioned image] Uses the whisper model to process audio transcription from classes and write summaries and following up questions 2023.06
@@
BOLLaMa 439 [Uncaptioned image] Implements an LLM interface to ease the usage of their BO code 2023.06
@@
text2concrete440 [Uncaptioned image] Uses ICL to predict compressive strength from concrete formulation 2023.06
MAPI_LLM441[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Database access and LLM prediction using ICL 2023.06
BO-LIFT442 [Uncaptioned image] Regression using ICL and text retrieval 2023.04
@@
ChemCrow45[Uncaptioned image] [Uncaptioned image] Molecular, cheminformatics, search and critique tools 2023.04
@@
It was Hocky and White 443 who discussed the early stages of models that could automate programming and, hence, the expected impacts in chemistry. Then, early work by White et al. 444 applied LLMs that could generate code to a benchmark set of chemical problems. In that case, not only were LLMs demonstrated to possess a notable understanding of chemistry, based on accurate question answering, but White et al. 444 imagined a potential to use them as base models to control knowledge augmentation and a variety of other tools. Thus, these LLMs could be used to execute routine tasks, optimize procedures, and enhance the retrieval of information from scientific literature across a range of scientific domains. To the best of our knowledge, this is the first review of autonomous agents in chemistry that have evolved since these two visionary conceptual perspectives. A deeper exploration follows below. One driving motivation for the need to augment LLMs with a more pertinent and dedicated knowledge base is the need to circumvent problems of a limited context prompt window, and the restriction that once an LLM is trained, any new information is beyond it’s reach since it necessarily has fallen outside its corpus of training data. Furthermore, LLMs are also known to hallucinate. Their predictions are probabilistic and, in science, if experimental evidence is available, then there is great value in building from known domain-specific information. Some improved prompt engineering can aid in the generation of results that are more likely to be accurate, but the use of autonomous agents may solve such problems completely in this next phase of AI in chemistry. In fact, even adding one or two components when building an agent, as opposed to a whole suite, has shown some significant gains.
@@
Building on this foundation, Ramos et al. 442 illustrated that LLMs could directly predict experimental outcomes from natural language descriptions, a technique they incorporated into a Bayesian optimization (BO) algorithm to streamline chemical processes. Because ICL was used, their approach did not require additional model training or finetuning, thereby greatly simplifying the optimization algorithm and removing the need to re-train or update models as data is gathered. Recently, Kristiadi et al. 445 demonstrated that similar results could also be achieved with a much smaller domain-specific model while performing parameter-efficient fine-tuning (PEFT) instead of ICL. Ranković and Schwaller 446 also explored BO using natural language. They used an LLM to encode chemical reaction procedures, described using natural language, and then trained a Gaussian process (GP) head to predict the reaction yield from the latent encoded representation of the procedure. This effective transfer-learning approach minimized training time by keeping the LLM frozen while updating only the MLP. Extending these ideas, Völker et al. 428 suggested sampling multiple model completions and adding a verifier model to select the next best procedure to continue the BO algorithm. They also used ICL and a short-term memory component to optimize alkali-activated concrete mix design. The authors thus underscored how augmenting with knowledge-driven design can outperform traditional data-driven design. These are examples of how agent-based approaches can execute complex optimization algorithms in a step-by-step approach, directly contributing to automation and leading to faster and more efficient experimental design.

### v2

@@ 4 LLM-based autonomous agents
The term “agent” originates in philosophy, referring to entities capable of making decisions371. Hence, in artificial intelligence, an “agent” is a system that can perceive its environment, make decisions, and act upon them in response to external stimuli372. Language has enabled humans to decide and act to make progress in response to the environment and its stimuli, and so LLMs are naturally ideal for serving as the core of autonomous agents. Thus, in agreement with Gao et al. 373, we define a “language agent” as a model or program (typically based on LLMs) that receives an observation from its environment and executes an action in this environment. Here, environment means a set of tools and a task. Hence, “LLM-based autonomous agents” refer to language agents whose core is based on an LLM model. Comprehensive analyses of these agents are available in the literature374, 372, 373, but this section highlights key aspects to prepare the reader for future discussions.
There is no agreed definition of the nomenclature to be used to discuss agents. For instance, Gao et al. 373 created a classification scheme that aims to group agents by their autonomy in biological research. This means a level 0 agent has no autonomy and can only be used as a tool, while a level 3 agent can independently create hypotheses, design experiments, and reason. Another previously suggested framework highlights agents’ construction, application, and evaluation as a parallel between agents and neural networks. Following this perspective, Wang et al. 374 categorizes agent components into four modules: profiling, memory, planning, and action. In contrast, Weng 375 also identifies four elements — memory, planning, action, and tools — but with a different emphasis. Meanwhile, Xi et al. 372 proposes a division into three components: brain, perception, and action, integrating profiling, memory, and planning within the brain component, where the brain is typically an LLM. Recently, Sumers et al. 376 proposed Cognitive Architectures for Language Agents (CoALA), a conceptual framework to generalize and ease the design of general-purpose cognitive language agents. In their framework, a larger cognitive architecture composed of modules and processes is defined. CoALA defines a memory, decision-making, and core processing module, in addition to an action space composed of both internal and external tools. While internal tools mainly interact with the memory to support decision-making, external tools compose the environment, as illustrated in Figure 5. Given a task that initiates the environment, the “decision process” runs continuously in a loop, receiving observations and executing actions until the task is completed. For more details, read Reference [376].
@@
In this review, we define an autonomous agent system as a model (typically an LLM) that continuously receives observations from the environment and executes actions to complete a provided task, as described by Gao et al. 373. Nevertheless, in contrast to CoALA373, we will rename “internal tools” as “agent modules” and “external tools” simply as “tools”, for clarity. The agent consists of trainable decision-making components such as the LLM itself, policy, memory, and reasoning scheme. In contrast, the environment comprises non-trainable elements like the task to be completed, Application Programming Interface (API) access, interfaces with self-driving labs, dataset access, and execution of external code. By referring to decision-making components as agent modules, we emphasize their inclusion as parts of the agent. By referring to non-trainable elements as tools, we highlight their role as part of the environment. We discuss six main types of actions. As shown in Figure 5, four of the six, memory, planning, reasoning, and profiling are agent modules. The remaining two actions (or tools) and perception are part of the environment. Since the perception is how the agent interacts with the environment and is not a trainable decision, we therefore included it as part of the environment.
@@ 4.1 Memory module
The role of the memory module is to store and recall information from past interactions and experiences to inform future decisions and actions. There are multiple types of memory in agents, namely sensory memory, short-term memory, and long-term memory. A major challenge in using agents is the limited context window, which restricts the amount of in-context information and can lead to information loss, thereby impacting the effectiveness of short-term and long-term memory. Solutions involve summarizing memory content377, compressing memories into vectors378, 379, 380, and utilizing vector databases381 or combinations thereof382, with various databases available such as ChromaDB, FAISS, Pinecone, Weaviate, Annoy, and ScaNN.383 Addressing these challenges to enhance agent memory continues to be a significant area of research384. Sensory, or procedural, memory is knowledge embedded into the model’s parameters during pretraining and/or in heuristics implemented into the agent’s code. Short-term, or working, memory includes the agent’s finite knowledge during a task, incorporating interaction history and techniques like in-context learning385 (ICL), which leverages the limited input’s context length for information retention. Long-term memory involves storing information externally, typically through an embedded vector representation in an external database. In the original CoALA373 paper, long-term memory is further divided into two different types of long-term memory: episodic, which registers previous experiences, and semantic, which stores general information about the world.
@@ 4.2 Planning and reasoning modules
The planning and reasoning module is made of two components. Planning involves identifying a sequence of actions required to achieve a specified goal. In the context of language agents, this means generating steps or strategies that the model can follow to solve a problem or answer a question, which can be enhanced with retrieval from previous experiences,386 and from feedback from post-execution reasoning387, 388. Reasoning refers to the process of drawing conclusions or making decisions based on available information and logical steps. For example, there are studies that demonstrate the benefits of LLM reasoning for question answering, where new context tokens can be integrated in a step-by-step way to guide the model towards more accurate answers389, 390, 391, 392, 393, 394. One popular reasoning strategy is Chain-of-Thought (CoT),391, 395, 396, 98, 397, 398 a reasoning strategy which substantially boosts QA performance by generating intermediate reasoning steps. CoT thereby reduces hallucinations and enhances interpretability, as demonstrated by improved results in models like PaLM399 and GPT-3 with benchmarks like GSM8K400, SVAMPs401, and MAWPS402.
@@
In advanced reasoning, final tasks are often decomposed into intermediary ones using a cascading approach, similar to Zero-shot-CoT390 and RePrompt387. However, while CoT is considered as single-path reasoning, CoT extensions like Tree-of-Thoughts393, Graph-of-Thoughts403, Self-consistent CoT392, and Algorithm-of-Thoughts404 offer multi-path reasoning. Furthermore, other models have pitted multiple agents against each other to debate or discuss various reasoning paths,405, 406, 407 while others use external planners to create plans408, 409. A feedback step during the execution of the plan was a further extension of the CoT ideas; this enables agents to refine their actions based on environmental responses adaptively, which is crucial for complex tasks.410, 411
@@
Another interesting reasoning scheme is the Chain-of-Verification(CoVe), 388 where once an answer is generated, another LLM is prompted to generate a set of verification questions to check for agreement between the original answer and the answers to the verification questions such that the final answer can be refined. The ReAct393 – Reason+Act – model proposes adding an observation step after acting. This means the LLM first reasons about the task and determines the necessary step for its execution, it performs the action and then observes the action’s result. Reasoning on that result, it can subsequently perform the following step. Similarly, Reflexion98 also implements a reasoning step after executing an action. However, Reflexion implements an evaluator and self-reflection LLMs to not only reason about each step but also to evaluate the current trajectory the agent is following using a long-term memory module. As the context increases, it may become challenging for agents to deal with the long prompt. Aiming to solve this issue, the Chain-of-Agents (CoA)412 extends reasoning schemes that leverage multi-agent collaboration to reason over long contexts. This framework employs workers and manager agents to process and synthesize information to generate the final response. CoA demonstrated improvements of up to 10% when compared against an RAG baseline.
@@
ReAct and Reflexion are closed-ended approaches where the agent starts with all the tools and must determine which to use. To address more open-world challenges, Wang et al. 413 introduced the Describe, Explain, Plan, and Select (DEPS) method, which extends this approach. Lastly, human inputs can also be used to provide feedback to the agent. Providing feedback using a human-in-the-loop approach is particularly interesting in fields where safety is a main concern.
@@ 4.3 Profiling module
LLMs can be configured to perform in specific roles, such as coders, professors, students, and domain experts, through a process known as profiling. Language agents can thus incorporate the profile through the LLM or through the agent code. The profiling approach involves inputting psychological characteristics to the agent, significantly impacting its decision-making process 414, 415, 416, 417. Profiling enables the creation of multi-agent systems that simulate societal interactions, with each agent embodying a unique persona within the group 386, 418. The most prevalent technique for profiling, called “handcrafting”, requires manually defining the agent’s profile, often through prompts or system messages 419, 420. While profiling can also be automated with LLMs 421, that automation method may only be suited for generating large numbers of agents since it offers less control over their overall behavior. An interesting application of profiling is the development of agent sets that reflect demographic distributions 422.
@@ 4.4 Perception
Perception is an analog to the human sensory system, which interprets multimodal information such as text, images, or auditory data, transforming it into a format comprehensible by LLMs, as demonstrated by SAM423, GPT4-V424, LLaVa425, Fuyu8B116, and BuboGPT426. In our proposed architecture, the perception is responsible for converting the task and the observations to a data representation that can be understood by the agent. Moreover, advancements in LLMs have led to the development of even more versatile models, such as the any-to-any Next-GPT117 and the any-to-text Macaw-LLM427. Employing multimodal LLMs, which are introduced above, in decision-making processes can simplify perception tasks for agents, with several studies exploring their use in autonomous systems428, 429.
@@ 4.5 Tools
The decision process is composed of three main steps: proposal, evaluation, and selection. During the proposal, one or more action candidates are selected using reasoning393, code structures430, 386, or simply by selecting every tool available431.389, 432, 392 The evaluation process consists of evaluating each selected action according to some metric to predict which action would bring more value to the agent. Lastly, the action is selected and executed.
Given that pretrained parameters (sensory memory) are limited, the model must use tools for complex tasks in order to provide reliable answers. However, LLMs need to learn how to interact with the action space and how and when to use those tools most accurately.433 LLMs can be pretrained or fine-tuned with examples of tool use, enabling them to operate tools and directly retrieve tool calls from sensory memory during a zero-shot generation.434 Recent studies investigate this approach, particularly focusing on open-source LLMs.435, 436, 437
As foundational AI models become more advanced, their abilities can be expanded. It was shown that general-purpose foundation models can reason and select tools even with no fine-tuning. For example, MRKL438 (pronounced “miracle”) implements an extendable set of specialized tools known as neuro-symbolic modules and a smart “router” system to retrieve the best module based on the textual input. These neuro-symbolic modules are designed to handle specific tasks or types of information, equipped with built-in capabilities and task-relevant knowledge. This pre-specialization allows the model to perform domain-specific tasks without needing a separate, domain-specific dataset. This design addresses the problem of LLMs lacking domain-specific knowledge and eliminates the need for the costly and time-consuming LLM fine-tuning step, using specialized data annotation.439 The router can receive support from a reasoning strategy to help select the tools439 or follow a previously created plan413. Recent advances have shown that LLMs can develop new tools of their own440, 441, 442, enabling agents to operate, as needed, in dynamic and unpredictable “open-worlds”, on unseen problems as illustrated by Voyager430. This capability allows agents to evolve and improve continually.
@@ 5 LLM-Based autonomous agents in scientific research
Table 4: Scientific LLM systems and agents. We identify the studies we classified as an agent with the icon [Uncaptioned image]443 and multi-agent systems with the icon [Uncaptioned image][Uncaptioned image]. [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]444 mean the agent bases his behavior on sensory, short, and long memory components, respectively. Besides the textual capabilities of LLM-based agents, [Uncaptioned image]445 and [Uncaptioned image]446 mean the agent has additional audio and visual perception, respectively.
@@
LLaMP447[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools for database access, literature search, and atomistic simulations 2024.06
SGA448[Uncaptioned image] [Uncaptioned image] Employ the LLMs in a optimization loop 2024.05
@@
CRISPR-GPT449[Uncaptioned image][Uncaptioned image] [Uncaptioned image] Tool for gene editing experiments design 2024.04
@@
TAIS450[Uncaptioned image][Uncaptioned image] [Uncaptioned image] Tools for gene expression data analysis 2024.02
@@
ChemReasoner451[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools for heuristic search, 3D structure generation, and prediction using GNNs 2024.02
SciAgent452[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Trained Mistral for tool usage. Evaluated it using MathToolBench’s tools 2024.02
STORM453 [Uncaptioned image][Uncaptioned image] Article writing using retrieval from multi-LLM conversations and pre-generated outline 2024.02
Völker et al. 454 [Uncaptioned image] Regression with ICL and text retrieval 2024.02
ProtAgent455 [Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools for proteins information retrieval, analyzing, de novo design, and 3D folded structure generation 2024.01
Organa370 [Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools for common lab procedures, reasoning about experimental results, and report writing 2024.01
PaperQA456[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools to search the scientific literature, gather evidence, and answer questions 2023.12
WikiCrow457[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Uses PaperQA as a tool 2023.12
Coscientist48[Uncaptioned image] [Uncaptioned image] Tools for running Python code, web-searching, and interacting with lab equipment 2023.12
Eunomia458[Uncaptioned image] [Uncaptioned image] Tools for literature and dataset searching and a chain-of–verification loop 2023.12
CALMS459[Uncaptioned image] [Uncaptioned image] Tools for using the Materials Project API, designing experiments, and using a hardware API to perform the experiment 2023.12
CoQuest460[Uncaptioned image] [Uncaptioned image] Research question generations and tools for literature visualization using a graph organization 2023.10
eXpertAI309[Uncaptioned image] [Uncaptioned image] Tools for applying XAI methods 2023.11
BioPlanner461[Uncaptioned image] [Uncaptioned image] Tool for protocol searching in the BioProt dataset 2023.10
IBM ChatChem462[Uncaptioned image] [Uncaptioned image] Tools for cheminformatics and accessing GT4SD and HuggingFace models 2023.09
ChatMOF463[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools for database search, property prediction, and MOF’s structure generation 2023.08
@@
AmadeusGPT47[Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools for writing and executing code for computer vision, machine learning, and spatial-temporal reasoning 2023.07
@@
i-Digest464 [Uncaptioned image] [Uncaptioned image] Uses the whisper model to process audio transcription from classes and write summaries and following up questions 2023.06
@@
BOLLaMa 465 [Uncaptioned image] Implements an LLM interface to ease the usage of their BO code 2023.06
@@
text2concrete466 [Uncaptioned image] Uses ICL to predict compressive strength from concrete formulation 2023.06
MAPI_LLM467[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Database access and LLM prediction using ICL 2023.06
BO-LIFT468 [Uncaptioned image] Regression using ICL and text retrieval 2023.04
@@
ChemCrow47[Uncaptioned image] [Uncaptioned image] Molecular, cheminformatics, search and critique tools 2023.04
@@
It was Hocky and White 469 who discussed the early stages of models that could automate programming and, hence, the expected impacts in chemistry. Then, early work by White et al. 470 applied LLMs that could generate code to a benchmark set of chemical problems. In that case, not only were LLMs demonstrated to possess a notable understanding of chemistry, based on accurate question answering, but White et al. 470 imagined a potential to use them as base models to control knowledge augmentation and a variety of other tools. Thus, these LLMs could be used to execute routine tasks, optimize procedures, and enhance the retrieval of information from scientific literature across a range of scientific domains. To the best of our knowledge, this is the first review of autonomous agents in chemistry that have evolved since these two visionary conceptual perspectives. A deeper exploration follows below. One driving motivation for the need to augment LLMs with a more pertinent and dedicated knowledge base is the need to circumvent problems of a limited context prompt window, and the restriction that once an LLM is trained, any new information is beyond it’s reach since it necessarily has fallen outside its corpus of training data. Furthermore, LLMs are also known to hallucinate. Their predictions are probabilistic and, in science, if experimental evidence is available, then there is great value in building from known domain-specific information. Some improved prompt engineering can aid in the generation of results that are more likely to be accurate, but the use of autonomous agents may solve such problems completely in this next phase of AI in chemistry. In fact, even adding one or two components when building an agent, as opposed to a whole suite, has shown some significant gains.
@@
Building on this foundation, Ramos et al. 468 illustrated that LLMs could directly predict experimental outcomes from natural language descriptions, a technique they incorporated into a Bayesian optimization (BO) algorithm to streamline chemical processes. Because ICL was used, their approach did not require additional model training or finetuning, thereby greatly simplifying the optimization algorithm and removing the need to re-train or update models as data is gathered. Recently, Kristiadi et al. 471 demonstrated that similar results could also be achieved with a much smaller domain-specific model while performing parameter-efficient fine-tuning (PEFT) instead of ICL. Ranković and Schwaller 472 also explored BO using natural language. They used an LLM to encode chemical reaction procedures, described using natural language, and then trained a Gaussian process (GP) head to predict the reaction yield from the latent encoded representation of the procedure. This effective transfer-learning approach minimized training time by keeping the LLM frozen while updating only the MLP. Extending these ideas, Völker et al. 454 suggested sampling multiple model completions and adding a verifier model to select the next best procedure to continue the BO algorithm. They also used ICL and a short-term memory component to optimize alkali-activated concrete mix design. The authors thus underscored how augmenting with knowledge-driven design can outperform traditional data-driven design. These are examples of how agent-based approaches can execute complex optimization algorithms in a step-by-step approach, directly contributing to automation and leading to faster and more efficient experimental design.